OpenStudy (anonymous):

Evaluate the polynomial x=9 and y=-5 x^2-3y^2+2 x y=

OpenStudy (anonymous):

put x=9 and y=-5 into x^2-3y^2+2 x y x^2-3y^2+2 x y=(9^2)-3(-5)^2+2(9)(-5) = -84